# From Grunts to Grammar: Emergent Language from Cooperative Foraging

### v1

@@ Abstract
Early cavemen relied on gestures, vocalizations, and simple signals to coordinate, plan, avoid predators, and share resources. Today, humans collaborate using complex languages to achieve remarkable results. What drives this evolution in communication? How does language emerge, adapt, and become vital for teamwork? Understanding the origins of language remains a challenge. A leading hypothesis in linguistics and anthropology posits that language evolved to meet the ecological and social demands of early human cooperation. Language did not arise in isolation, but through shared survival goals. Inspired by this view, we investigate the emergence of language in multi-agent Foraging Games. These environments are designed to reflect the cognitive and ecological constraints believed to have influenced the evolution of communication. Agents operate in a shared grid world with only partial knowledge about other agents and the environment, and must coordinate to complete games like picking up high-value targets or executing temporally ordered actions. Using end-to-end deep reinforcement learning, agents learn both actions and communication strategies from scratch. We find that agents develop communication protocols with hallmark features of natural language: arbitrariness, interchangeability, displacement, cultural transmission, and compositionality. We quantify each property and analyze how different factors, such as population size and temporal dependencies, shape specific aspects of the emergent language. Our framework serves as a platform for studying how language can evolve from partial observability, temporal reasoning, and cooperative goals in embodied multi-agent settings. We will release all data, code, and models publicly.
@@
Figure 1: An Overview of Foraging Games. (a) Two agents operate in a $5\times 5$ partially observable grid world. Each agent can observe a $3\times 3$ grid centered on itself. The two agents are required to pick up the goal items simultaneously. (b) ScoreG game: Two items are assigned random scores; each agent observes only one score, and both must pick up the higher-scoring item to succeed in an episode. This game is designed to encourage agents to communicate about items’ scores. (c) TemporalG game: Two items are assigned random spawn times; each agent observes only one item when it spawns. The agents must pick up the items in the same order as their spawn times—i.e., the earlier item must be picked up first. This game is designed to encourage agents to communicate about time.
1 Introduction
Before the emergence of structured language, early humans relied on gestures, vocalizations, and simple signals to coordinate their actions in a partially observable and socially interdependent world. A leading hypothesis is that human language evolved not in isolation, but through cooperative interaction under partial observability sterelny2012language ; tomasello2010origins ; nowak2000evolution ; christiansen2003language ; nowak2001towards . Language is not merely an abstract code but a tool shaped by social use and shared goals wittgenstein2009philosophical ; wagner2003progress . Although direct evidence from these early stages is unavailable, multi-agent simulations offer a way to study how language emerges from the need to coordinate with others cangelosi2012simulating ; kirby2014iterated ; lazaridou2017multi ; lazaridou2018emergence .
Prior work on emergent communication has primarily focused on referential games (RG), where a speaker conveys task-relevant information to a listener over a limited communication channel lazaridou2017multi ; lewis2008convention ; kharitonov2019egg ; gualdoni2024bridging . These settings have advanced our understanding of symbol grounding and compositionality, but often impose simplifying assumptions: agents are limited to unidirectional communication galke2022emergent , operate in fixed speaker-listener roles, and perform disembodied tasks with passive input processing and no active interaction with the environment. Even recent extensions to population-based learning kim2021emergent ; chaabouni2022emergent ; michelrevisiting typically decouple communication from physical behavior. Such setups diverge from the ecologically grounded, socially interdependent, and embodied conditions under which human language likely evolved tomasello2010origins ; dessalles2007we .
To bridge this gap, we introduce Foraging Games (FG), a multi-agent framework designed to reflect more realistic ecological and cognitive constraints that may have shaped early language dessalles2007we ; tomasello2010origins . FG supports bidirectional communication. Agents are trained to jointly learn both physical action and communication, as early humans likely did during cooperative foraging. The environment enforces embodiment: agents must explore, observe, and act within a dynamic and partially observable world that they can influence. Moreover, each agent interacts with a population of diverse partners, facilitating the study of generalization to new partners, dialect formation, and cultural transmission kim2021emergent ; rita2022role ; michelrevisiting .
Specifically, we formulate FG as a partially observable grid world in which agents must complete multi-step games as shown in Figure 1. Success depends not only on taking effective actions, but also on communicating what they see and know. It includes two games: (1) collecting the most valuable item with a partner, which encourages communication about items’ scores and locations, and (2) picking up items in a specific order with a partner, which requires communication about when items were seen. Each agent has deep neural network modules for perception, memory, and policies to translate internal representations into actions and messages. Agents communicate through discrete messages drawn from a finite set of learnable vocabulary exchanged at each time step. They are trained independently using Proximal Policy Optimization (PPO) schulman2017proximal , with no shared modules, parameters, or gradients, reflecting the personalized and decentralized nature of each agent.
@@ 1 Introduction
Subsequently, agents achieve a success rate (SR) above $95\%$ across all games. Although two trained agents perform well when paired with each other, each fails when paired with a copy of itself at test time, suggesting that an agent understands only its partner’s language, not the one it produces. We propose two solutions to this problem. The first is to train a population with more than two agents. We hypothesize that developing a shared language becomes the optimal strategy when agents must communicate with multiple partners. The second is to incorporate self-interaction during training charakorn2023generating ; charakorn2024diversity , motivated by the observation that humans can speak to themselves. These solutions encourage agents to comprehend their own messages, promoting convergence on a shared common language and reflecting the property of interchangeability. Since human language develops in populations, we further explore language properties in groups of agents connected via fully connected and ring-structured social networks. We decode task-relevant information, such as item positions, from the agents’ messages using logistic regression, verifying that the messages are meaningful rather than random. Above-chance decoding accuracy indicates that a language has emerged to communicate item properties. Finally, we show that agents can develop implicit communication wang2025scout ; grupen2022multi ; dreyer2025comparing , i.e., gaining information by observing a partner’s behavior, when they are unable to send messages. Our key contributions are as follows:
@@
1. We introduce Foraging Games, a framework for studying emergent communication in physically grounded, socially interdependent environments. FG provides ecological and cognitive constraints that early humans might have faced, including embodiment, behavior coordination, partial observability, bidirectional communication, and temporal reasoning.
2. We introduce a hybrid cross-and-self-play training regime that encourages interchangeable language and supports cultural transmission: agents that are closer in the training population develop more similar languages than those farther apart. While agents need not converge on a single shared language, reflecting the arbitrariness of language, their messages remain mutually intelligible within their social context. As an ablation, we observe that without this regime, e.g., in purely decentralized cross-play training, two trained agents often develop non-interchangeable languages and even fail to understand their own messages when roles are swapped at test time.
3. We design a series of tasks in FG that elicit both temporal and spatial displacement in emergent language. Agents develop time adverbials lipinski2023s that refer to when past events occurred, along with messages indicating the location of those events. Rigorous analysis of messages from trained agents across various social networks enables systematic and quantitative evaluation of compositionality and other key attributes characteristic of emergent communication.
@@ 2 Related Work
Emergent communication (EC) provides insight into how human language may have evolved lazaridou2017multi ; lazaridou2018emergence ; kirby2014iterated ; chaabouni2022emergent ; cangelosi1998emergence ; cangelosi2001evolution ; boldtreview ; peters2025emergent ; lazaridou2020emergent , as well as how to develop more effective representations mu2021emergent ; carmelictd ; zionsemantics ; denamganai2024erelela ; yao2022linking . Early work in this field relied on evolutionary computation cangelosi1998emergence ; cangelosi2001evolution or Bayesian modeling nowak2000evolution ; kirby2007innateness to simulate language emergence and its underlying dynamics. With the advent of deep reinforcement learning (DRL) mnih2013playing ; mnih2015human , a more powerful and flexible computational framework has emerged for studying EC in realistic and complex settings foerster2016learning . Modern EC studies apply DRL in Lewis-style referential games lewis2008convention , where a speaker must convey task-relevant messages to a listener through a limited-capacity communication channel. The game’s objective typically involves either target identification or input reconstruction zionsemantics ; gualdoni2024bridging ; lipinski2024speaking ; kharitonov2019egg . Since language is inherently a social construct developed by communities rather than individuals, recent work has shifted toward examining the properties of language in populations of speakers and listeners kim2021emergent ; chaabouni2022emergent ; rita2022role ; michelrevisiting . The resulting frameworks often resemble variational autoencoders (VAEs) kingma2013auto ; higgins2017beta ; uedalewis , where language acts as a compressed latent code, rather than a flexible medium for coordination in the world.
Unlike most prior setups, our agents are bidirectional communicators evtimova2018emergent ; taillandier2023neural ; graesser2019emergent ; kottur2017natural ; nikolaus2023emergent and embodied in the environment jain2019two ; patel2021interpretation ; mu2023ec2 . This setup better reflects real-world communication, where agents must both produce and interpret signals to collaboratively complete a task effectively. Crucially, communication and sequential decision-making co-evolve: agents learn what to say as they learn how to act mordatch2018emergence ; kajic2020learning . They must actively explore, observe, remember, and adapt to their partners’ behavior to succeed, resulting in dynamic, context-dependent coordination.
The closest work to ours is mordatch2018emergence , which also studies emergent communication in a physically interactive setting. However, their agents do not coordinate actions to achieve a shared goal, and all agents share a single deep neural network. Moreover, they rely on a differentiable environment that allows gradients to flow directly from the environment to the neural network. In contrast, our setting is fully decentralized and reward-driven, where communication and coordination emerge solely from reward, thus emulating the constraints and limitations faced by early humans.
We position our work as an initial step toward studying language emergence within a novel embodied framework, one in which language is not an isolated code but a survival tool. Within FG, agents must continually integrate perception, memory, and interaction to align their actions, mirroring the ecological and social constraints believed to have shaped human language.
@@
Figure 2: A graphical overview of our method. (a) Agents exchange messages at every time step. (b) The neural architecture of a single agent. On the input side, the received integer message is mapped to a real-valued vector via a learnable lookup table and passed through the message encoder. The grid observation and agent position are processed by the grid encoder and position encoder, respectively. The outputs of all three encoders are concatenated and passed to an LSTM module, which maintains temporal memory. On the output side, the message head, action head, and value head produce the next message token, environment action, and estimated state value (See Sec. 4.1).
@@ 3 Foraging Games (FG)
Foraging-style games are used as benchmarks for multi-agent reinforcement learning (MARL) algorithms albrecht2013gamefg ; papoudakis2021agentfg ; yang2022transformerfg ; jafferjee2022tamingfg . However, existing environments often assume full observability and are not designed to support the study of emergent communication in humans. We develop our FG, in which decentralized agents share a common goal and must navigate to and simultaneously pick up the goal items to earn equal rewards. Our FG is designed to encourage agents to communicate about what they observed. Each agent is provided with only partial knowledge of the environment (e.g., observing only one item). Agents must reach a consensus based on their perspectives by communicating the message from a fixed-size dictionary with learnable embeddings and choosing appropriate actions, where the action space is {move left, move right, move up, move down, and pick up}. FG has two games: ScoreG and TemporalG. The former is used to investigate four properties of emergent language: interchangeability (agents can understand the messages they produce), arbitrariness (symbols acquire meaning through social agreement rather than inherent structure), compositionality (messages are built from reusable parts reflecting task semantics), and cultural transmission (language patterns are passed across agents), under varying population sizes and training regimes. The latter focuses on displacement, aiming to determine whether agents’ messages encode temporal and spatial information about seen items in the past.
@@
is set in a $2D$ grid world with the size of $5\times 5$, containing $2$ items, $2$ agents, and a wall, as shown in Figure 1. The wall refers to the area outside the $5\times 5$ grid. Each grid cell can be occupied by either an agent or an item, but not both simultaneously. Agents cannot move to the wall, the occupied grid cell, or the same grid cell. Each agent receives an observation $\mathbf{o}^{(t)}=(\mathbf{x}^{(t)},\mathbf{p}^{(t)})$, where $\mathbf{x}^{(t)}\in\mathbb{R}^{3\times 3\times 2}$ ($3$ is the receptive field size and 2 refers to two channels). The first channel of the receptive field $\mathbf{x}^{(t)}$ represents an occupancy map: a value of zero indicates that a grid cell contains neither an item nor a wall. The second channel shows the item’s score. The receptive field of each agent has its center at the agent and observes the agent’s surroundings, delineated as a red box in Figure 1. To simulate incomplete knowledge, each agent observes the score of one item during an episode. The agent’s absolute position in the environment is given by $\mathbf{p}^{(t)}\in\mathbb{R}^{2}$. The goal item is defined as the item with the highest score. An episode is deemed successful if both agents navigate to and simultaneously pick up the goal item. In this case, they receive a shared positive reward of $+1$; otherwise, they receive a shared negative reward of $-1$. Additionally, similar to wijmansdd , successful episodes grant both agents a bonus reward $r_{b}$ based on their efficiency: $r_{b}=\frac{T_{\max}-T_{\text{total}}}{T_{\max}}$, where $T_{\max}=10$ is the maximum number of steps allowed per episode, and $T_{\text{total}}$ is the total number of steps taken by the agents. The bonus is to encourage the agents to complete the game as soon as possible. To prevent agents from overfitting to specific scores in the ScoreG task while still evaluating in-distribution generalization, we use a held-out set of test scores that do not overlap with those seen during training. Specifically, training scores are randomly sampled from the set $\{5,10,15,\dots,250\}$, while test scores are sampled from the disjoint set $\{2,4,6,8,12,\dots,248\}$. This setup ensures that test scores lie within the same distributional range as the training scores, while remaining unseen during training. To isolate the role of explicit communication, agents are made invisible to one another by default, i.e., a grid cell occupied by another agent appears empty. This design prevents any form of implicit communication through visual cues wang2025scout ; grupen2022multi ; li2021deep ; li2024explicit .
@@
Two agents must cooperatively pick up two items by first navigating to and simultaneously collecting the first spawned item, followed by the second. The pickup must occur in the same order as the items appear. The environment follows the same configuration as ScoreG, except the agent’s observation $\mathbf{x}^{(t)}$ excludes the score channel, and the episode length is limited to $T_{\text{max}}=20$. Initially, agents are spawned on opposite sides of the grid and remain frozen at time step $t=1$ to $t=6$. Subsequently, each item appears at a distinct time step, drawn uniformly from the fixed duration set $\{1,2,3,\ldots,6\}$. Each item is guaranteed to appear within at least one agent’s receptive field, ensuring it is visible when it spawns. Agents begin moving and attempting to pick up the items only after $t=6$. To promote both spatial and temporal displacement, we restrict the communication range so that agents can only successfully send messages if they occupy adjacent grid cells. Otherwise, a zero-value message is sent. Reward is given in the same way as in ScoreG.

### v2

# From Grunts to Lexicons: Emergent Language from Cooperative Foraging

@@ Abstract
Language is a powerful communicative and cognitive tool. It enables humans to express thoughts, share intentions, and reason about complex phenomena. Despite our fluency in using and understanding language, the question of how it arises and evolves over time remains unsolved. A leading hypothesis in linguistics and anthropology posits that language evolved to meet the ecological and social demands of early human cooperation. Language did not arise in isolation, but through shared survival goals. Inspired by this view, we investigate the emergence of language in multi-agent Foraging Games. These environments are designed to reflect the cognitive and ecological constraints believed to have influenced the evolution of communication. Agents operate in a shared grid world with only partial knowledge about other agents and the environment, and must coordinate to complete games like picking up high-value targets or executing temporally ordered actions. Using end-to-end deep reinforcement learning, agents learn both actions and communication strategies from scratch. We find that agents develop communication protocols with hallmark features of natural language: arbitrariness, interchangeability, displacement, cultural transmission, and compositionality. We quantify each property and analyze how different factors, such as population size, social dynamics, and temporal dependencies, shape specific aspects of the emergent language. Our framework serves as a platform for studying how language can evolve from partial observability, temporal reasoning, and cooperative goals in embodied multi-agent settings. We will release all data, code, and models publicly.
1 Introduction
The evolution of human language remains a central open question in science. While humans fluently use and understand language, its origins are still debated. A prominent hypothesis suggests that language emerged through cooperative interaction under partial observability sterelny2012language ; tomasello2010origins ; nowak2000evolution ; christiansen2003language ; nowak2001towards . Rather than an abstract code, language is viewed as a tool shaped by social use and shared goals wittgenstein2009philosophical ; wagner2003progress . Although direct evidence from early stages is lacking, multi-agent simulations provide a means to investigate how language might arise from the need to coordinate with others cangelosi2012simulating ; kirby2014iterated ; lazaridou2017multi ; lazaridou2018emergence .
Prior work on Emergent Communication (EC) has primarily focused on referential games (RG), where a speaker conveys task-relevant information to a listener over a limited communication channel lazaridou2017multi ; lewis2008convention ; kharitonov2019egg ; gualdoni2024bridging . These settings have advanced our understanding of symbol grounding and compositionality, but often impose simplifying assumptions: agents are limited to unidirectional communication galke2022emergent , operate in fixed speaker-listener roles, and perform disembodied tasks with passive input processing and no active interaction with the environment. Even recent extensions to population-based learning dubova2020effects ; dubova2020reinforcement ; kim2021emergent ; chaabouni2022emergent ; michelrevisiting typically decouple communication from physical behavior. Such setups diverge from the ecologically grounded, socially interdependent, and embodied conditions under which human language likely evolved tomasello2010origins ; dessalles2007we .
To bridge this gap, we introduce Foraging Games (FG), a multi-agent framework designed to reflect more realistic ecological and cognitive constraints that may have shaped early human language dessalles2007we ; tomasello2010origins . FG supports bidirectional communication. Agents are trained to jointly learn both physical action and communication, as early humans likely did during cooperative foraging. The environment enforces embodiment: agents must explore, observe, and act within a dynamic and partially observable world that they can influence. Moreover, each agent interacts with a population of diverse partners, facilitating the study of generalization to new partners, dialect formation, and cultural transmission kim2021emergent ; rita2022role ; michelrevisiting .
Specifically, we formulate FG as a partially observable grid world in which agents must complete multi-step games as shown in Figure 1. Success depends not only on taking effective actions, but also on communicating what they see and know. It includes two games: (1) collecting the most valuable item with a partner, which encourages communication about items’ scores and locations, and (2) picking up items in a specific order with a partner, which requires communication about when items were seen. Each agent has deep neural network modules for perception, memory, and policies to translate internal representations into actions and messages. Agents communicate through discrete messages drawn from a finite set of learnable vocabulary exchanged at each time step. They are trained independently using Proximal Policy Optimization (PPO) schulman2017proximal , with no shared parameters, or gradients, reflecting the personalized and decentralized nature among agents.
@@ 1 Introduction
Figure 1: An Overview of Foraging Games. (a) Two agents operate in a $5\times 5$ partially observable grid world. Each agent can observe a $3\times 3$ grid centered on itself. The two agents are required to pick up the goal items simultaneously. (b) ScoreG game: Two items are assigned random scores; each agent observes only one score, and both must pick up the higher-scoring item to succeed in an episode. This game is designed to encourage agents to communicate about items’ scores. (c) TemporalG game: Two items spawn at random times, each observed by only one agent. Agents must collect them in chronological order, encouraging communication about time.
With this framework in place, we aim to provide answers to the following questions: (Q1) What properties does emergent communication have when it arises in embodied, interactive, and collaborative agents? (Q2) Do agents truly understand the language they use, and what drives the emergence of a shared language when the same agent acts as both speaker and listener? (Q3) How do population size and social dynamics affect the emergent language? (Q4) Can agents refer to spatial and temporal events through their language? (Q5) Can agents communicate implicitly through non-verbal behaviors, such as body movements?
@@
Empirically, we find that agents achieve a success rate (SR) above $95\%$ across all games. Although two trained agents perform well when paired with each other, each fails when paired with a copy of itself at test time, suggesting that an agent understands only its partner’s language, not the one it produces. We propose two solutions to this problem. The first is to train a population with more than two agents. We hypothesize that developing a shared language becomes the optimal strategy when agents must communicate with multiple partners. The second is to incorporate self-interaction during training dubova2020effects ; charakorn2023generating ; charakorn2024diversity , motivated by the observation that humans can speak to themselves. These solutions encourage agents to comprehend their own messages, promoting convergence on a shared common language and reflecting the property of interchangeability. Since human language develops in populations, we further explore language properties in groups of agents connected via different social structures: fully-connected, ring-structured, and small-world-structured networks. We decode task-relevant information, such as items’ positions, scores, and spawn times, from the agents’ messages using logistic regression, verifying that the messages are meaningful rather than random. Above-chance decoding accuracy indicates that a language has emerged to communicate item properties. Specifically, agents develop time adverbials lipinski2023s that refer to when past events occurred, along with messages indicating the location of those events. Finally, we show that agents can develop implicit communication wang2025scout ; grupen2022multi ; dreyer2025comparing , i.e., gaining information by observing a partner’s behavior, when they are unable to send messages.
Our novel contributions are as follows:
@@
1. We introduce Foraging Games, a framework for studying emergent communication that provides ecological and cognitive constraints resembling those faced by early humans, including embodiment, behavior coordination, partial observability, bidirectional communication, and temporal reasoning.
2. We propose a hybrid cross-and-self-play training regime that enhances cultural transmission in weakly-connected social networks: agents that are closer in the training population develop more similar languages than those farther apart.
3. We design a series of tasks in FG that elicit both temporal and spatial displacement in emergent language, which could not arise without embodiment.
@@ 2 Related Work
EC offers insight into the evolution of human language lazaridou2017multi ; lazaridou2018emergence ; kirby2014iterated ; chaabouni2022emergent ; cangelosi1998emergence ; cangelosi2001evolution ; boldtreview ; peters2025emergent ; lazaridou2020emergent and the development of more effective representations mu2021emergent ; carmelictd ; zionsemantics ; denamganai2024erelela ; yao2022linking . Early studies employed evolutionary computation cangelosi1998emergence ; cangelosi2001evolution or Bayesian modeling nowak2000evolution ; kirby2007innateness to simulate language emergence and its dynamics. The advent of deep reinforcement learning (DRL) mnih2013playing ; mnih2015human introduced a more powerful framework for studying EC in realistic, complex settings foerster2016learning . Contemporary work often applies DRL to Lewis-style referential games (RG) lewis2008convention , where a speaker conveys task-relevant information to a listener through a limited communication channel, typically for target identification or input reconstruction zionsemantics ; gualdoni2024bridging ; lipinski2024speaking ; kharitonov2019egg . Because language is inherently social, research has shifted toward population-level studies dubova2020effects ; dubova2020reinforcement ; kim2021emergent ; chaabouni2022emergent ; rita2022role ; michelrevisiting , examining how social dynamics and population training influence consistency dubova2020effects ; dubova2020reinforcement and dialect formation kim2021emergent . However, many resulting frameworks resemble autoencoders kingma2013auto ; higgins2017beta ; uedalewis , treating language as a compressed latent code rather than a flexible medium for real-world coordination.
Recent studies have examined more realistic conditions, such as bidirectional exchange, where agents both send and receive messages evtimova2018emergent ; dubova2020effects ; dubova2020reinforcement ; taillandier2023neural ; graesser2019emergent ; kottur2017natural ; nikolaus2023emergent , and embodiment jain2019two ; patel2021interpretation ; mordatch2018emergence ; kajic2020learning , where agents learn to communicate through interaction with their environment and partners. Only two studies mordatch2018emergence ; jain2019two incorporate both. Most embodied EC work jain2019two ; patel2021interpretation prioritizes task performance over analyzing EC itself and omits biologically plausible decentralized training. Another line of research uses EC as a control prompt, outperforming natural language in embodied tasks mu2023ec2 , though it is still learned through disembodied RG. The closest work to ours is mordatch2018emergence , which investigates EC in a physical setting but with all agents sharing a single network, lacking coordinated action toward a shared goal. It also overlooks population training, social dynamics, and key properties such as compositionality, interchangeability, and displacement, while relying on a differentiable environment with direct gradient flow. By contrast, our approach is fully decentralized and reward-driven, with communication and coordination emerging solely from reward, reflecting constraints faced by early humans. We further demonstrate that simple FG yields core linguistic phenomena—most notably displacement, grounded in embodiment.
@@
Figure 2: A graphical overview of our method. (a) Agents exchange messages at every time step. (b) The neural architecture of a single agent. On the input side, the received integer message is mapped to a real-valued vector via a learnable lookup table and passed through the message encoder. The grid observation and agent position are processed by the grid encoder and position encoder, respectively. The outputs of all three encoders are concatenated and passed to an LSTM module, which maintains temporal memory. On the output side, the message head, action head, and value head produce the next message token, environment action, and estimated state value.
@@ 3 Foraging Games (FG)
Foraging-style games are used as benchmarks for multi-agent reinforcement learning (MARL) algorithms albrecht2013gamefg ; papoudakis2021agentfg ; yang2022transformerfg ; jafferjee2022tamingfg ; jaques2019social ; ikram2021hexajungle . However, existing environments often assume full observability and are not designed to support the study of emergent communication. We develop our FG, in which decentralized agents share a common goal and must navigate to and simultaneously pick up the goal items to earn equal rewards. Our FG is designed to encourage agents to communicate bidirectionally about their observations. Each agent is provided with only partial knowledge of the environment (e.g., observing only one item). Agents must reach a consensus based on their perspectives by communicating the message from a fixed-size dictionary with learnable embeddings and choosing appropriate actions, where the action space is {move left, move right, move up, move down, and pick up}. FG has two games: ScoreG and TemporalG. The former is used to investigate four properties of emergent language: interchangeability (agents can understand the messages they produce), arbitrariness (symbols acquire meaning through social agreement rather than inherent structure), compositionality (messages are built from reusable parts reflecting task semantics), and cultural transmission (language patterns are passed across agents), under varying population sizes and training regimes. The latter focuses on displacement, aiming to determine whether agents’ messages encode temporal and spatial information about seen items in the past.
@@
is set in a $2D$ grid world with the size of $5\times 5$, containing $2$ items, $2$ agents, and a wall, as shown in Figure 1. The wall refers to the area outside the $5\times 5$ grid. Each grid cell can be occupied by either an agent or an item, but not both simultaneously. Agents cannot move to the wall, the occupied grid cell, or the same grid cell. Each agent receives an observation $\mathbf{o}^{(t)}=(\mathbf{x}^{(t)},\mathbf{p}^{(t)})$, where $\mathbf{x}^{(t)}\in\mathbb{R}^{3\times 3\times 2}$ (33 is the receptive field size and 2 refers to two channels). The first channel of the receptive field $\mathbf{x}^{(t)}$ represents an occupancy map: a value of zero indicates that a grid cell contains neither an item nor a wall. The second channel shows the item’s score. The receptive field of each agent has its center at the agent and observes the agent’s surroundings, delineated as a red box in Figure 1. To simulate incomplete knowledge, each agent observes the score of one item during an episode. The agent’s absolute position in the environment is given by $\mathbf{p}^{(t)}\in\mathbb{R}^{2}$. The goal item is defined as the item with the highest score. An episode is deemed successful if both agents navigate to and simultaneously pick up the goal item. In this case, they receive a shared positive reward of +1+1; otherwise, they receive a shared negative reward of 1-1. Additionally, similar to wijmansdd , successful episodes grant both agents a bonus reward $r_{b}$ based on their efficiency: $r_{b}=\frac{T_{\max}-T_{\text{total}}}{T_{\max}}$, where $T_{\max}=10$ is the maximum number of steps allowed per episode, and $T_{\text{total}}$ is the total number of steps taken by the agents. The bonus is to encourage the agents to complete the game as soon as possible. To prevent agents from overfitting to specific scores in the ScoreG task while still evaluating in-distribution generalization, we use a held-out set of test scores that do not overlap with those seen during training. Specifically, training scores are randomly sampled from the set $\{5,10,15,\dots,250\}$, while test scores are sampled from the disjoint set $\{2,4,6,8,12,\dots,248\}$. This setup ensures that test scores lie within the same distributional range as the training scores, while remaining unseen during training. We also include the game variant to unseen item positions in Sec. D.5. To isolate the role of explicit communication, agents are made invisible to one another by default, i.e., a grid cell occupied by another agent appears empty. This design prevents any form of implicit communication through visual cues wang2025scout ; grupen2022multi ; li2021deep ; li2024explicit .
@@
Two agents must cooperatively pick up two items by first navigating to and simultaneously collecting the first spawned item, followed by the second. The pickup must occur in the same order as the items appear. The environment follows the same configuration as ScoreG, except the agent’s observation $\mathbf{x}^{(t)}$ excludes the score channel, and the episode length is limited to $T_{\text{max}}=20$. Initially, agents are spawned on opposite sides of the grid and remain frozen at time step $t=1$ to $t=6$. Subsequently, each item appears at a distinct time step, drawn uniformly from the fixed duration set $\{1,2,3,\ldots,6\}$. Each item is guaranteed to appear within at least one agent’s receptive field, ensuring it is visible when it spawns. Agents begin moving and attempting to pick up the items only after $t=6$. To encourage both spatial and temporal displacement, communication is restricted to adjacent grid cells; otherwise, a zero-value message is transmitted. Rewards follow the same scheme as in ScoreG.